Taylorj33:

The circumference of a circle is 5π ft. What is the area, in square feet? Express your answer in terms of piπ.

sxlvergxge:

Please ask this question in another form! People aren't supposed to solve the problem for you! Perhaps ask a friend, teacher, or family member for help? If you have notes on the subject you might want to look for those! For now we can't do much! Good luck!

surjithayer:

\[circumference~of~circle=2\pi r\] \[area~of~circle=\pi r^2\]

curriful:

The circumference of a circle is given by the formula c=2 pi r, where r is the radius. In this problem, the circumference is given as 5pi ft. By rearranging the formula, we can find the radius, r=C/(2pi). Substituting the given value the radius is 5pi ft / (2pi) = 5/2 ft. The area of a circle is given by the formula A = pi r^2. Substituting the value of the radius, we get A = pi (5/2 ft)^2 = 25/4 pi ft^2. Therefore, the area of the circle is 25/4pi square feet.

curriful:

@curriful wrote:
The circumference of a circle is given by the formula c=2 pi r, where r is the radius. In this problem, the circumference is given as 5pi ft. By rearranging the formula, we can find the radius, r=C/(2pi). Substituting the given value the radius is 5pi ft / (2pi) = 5/2 ft. The area of a circle is given by the formula A = pi r^2. Substituting the value of the radius, we get A = pi (5/2 ft)^2 = 25/4 pi ft^2. Therefore, the area of the circle is 25/4pi square feet.
i dont know if i did this right, is it right? -

surjithayer:

\[2\pi r=5\pi ,r=\frac{ 5\pi }{ 2\pi}=\frac{ 5 }{ 2 }\] \[area~A=\pi r^2=\pi(\frac{ 5 }{ 2 })^2=\frac{ 25 }{ 4}~sq.~ft.\]

curriful:

@surjithayer wrote:
\[2\pi r=5\pi ,r=\frac{ 5\pi }{ 2\pi}=\frac{ 5 }{ 2 }\] \[area~A=\pi r^2=\pi(\frac{ 5 }{ 2 })^2=\frac{ 25 }{ 4}~sq.~ft.\]
thanks for this, i wasnt sure if i was right or not -